Hi,
Am trying to help my autistic boy and unfortunately am not very good with math. He has a Age Word Problem as below:

Bill and Ben are brothers the sum of their ages is equal to 4 times Bob's age who is half as old as Bridget. The average age of the 4 children is 8. Write this as an equation, then solve it.
I tried to work out the equations and came out with something as below:
A=Bill's age
B=Ben's age
C=Bob's age
D=Bridget's age
A + B = 4C
D = 2C
A + B + C + D = 32
When I tried to work out the values, am not coming up with a whole number as the answer. Also am getting an invalid age value for A and B as well. What am I doing wrong?
Any advice will be much appreciated. Thanks in advance.


Answer by Alan3354(69443)

Bill and Ben are brothers the sum of their ages is equal to 4 times Bob's age who is half as old as Bridget. The average age of the 4 children is 8. Write this as an equation, then solve it.
I tried to work out the equations and came out with something as below:
A=Bill's age
B=Ben's age
C=Bob's age
D=Bridget's age
A + B = 4C
D = 2C
A + B + C + D = 32
When I tried to work out the values, am not coming up with a whole number as the answer. Also am getting an invalid age value for A and B as well. What am I doing wrong?
---------------
A + B + C + D = 32
4C + C + 2C = 32
C = 32/7
-------
D = 64/7
-----
A + B = 128/7
A & B cannot be determined with the data given, only the sum.
-----------------
You would expect integer solutions, but they're not.
Maybe a typo somewhere.